Solve the given differential equation by variation of parameters. 2x²y + 3xy - y xi
Answer #1

OS - 2x2y+3x4-y = x35x solution to homogeneas port 21²4 tany-y=0 (40) = untczaka particular solution using Varaten of parY1 g(x) du W(0) * x3ja da og 2 x 3/2 s x32.x²Jse da = 2 // Jarda 2 3 PS = ? as IS How, p=4291+4242 = 7 2 2 + 2 x 5 x 2 39 کر)